OpenStudy (anonymous):

the product of two whole numbers is 36. What is the greatest possible sum the number can have?

OpenStudy (anonymous):

The divisors of 36 are 1, 2, 3, 4, 6, 9, 12, 18, and 36. We can see that when paired properly the pair with the greatest sum is 36+1=37

OpenStudy (anonymous):



OpenStudy (anonymous):

@dacelin_naomie

myininaya (myininaya):

xy=36 S=x+y We want to maximize S, correct?

OpenStudy (ybarrap):

We are given that \(a\times b=36\) and need to maximize \(a+b\). Applying The Method of La Grange Multipliers: $$ g(a,b,\lambda)=a+b+\lambda(a\times b-36)\\ g_a=1+\lambda a=0\implies a=\cfrac{-1}{\lambda}\\ g_b=1+\lambda b=0\implies b=\cfrac{-1}{\lambda}\\ g_\lambda=a\times b-36\implies \cfrac{1}{\lambda^2}=36\\ \lambda=\pm\cfrac{1}{6}\\ a=\pm6\\ b=\pm6\\ $$ So, the greatest possible sum the number can have is \(6+6=12\).

OpenStudy (anonymous):

hey @myininaya am i correct?

OpenStudy (anonymous):

@ybarrap 36+1=37 ----------------------36*1=36

myininaya (myininaya):

Hmmm...I think you minimized instead of maximized @ybarrap

myininaya (myininaya):

And yes @Uditkulka looks good.

OpenStudy (ybarrap):

I agree, looks like I minimized. @Uditkulka 's solution looks right.

myininaya (myininaya):

I wanted to do some calculus too @ybarrap :(